Find the point on the y-axis that is equidistant from (-5,-2) and (3,2)

Thank You!!!

Answer by stanbon(75887)

Find the point on the y-axis that is equidistant from (-5,-2) and (3,2)
----------
Let the point be (0,y)
-----
Equation:
distance from (-5,-2) = distance from (3,2)
----
(y--2)^2 + (0--5)^2 = (y-2)^2 + (0-3)^2
------
(y+2)^2 + 25 = (y-2)^2 + 9
------
y^2 + 4y + 29 = y^2-4y+13
------
8y = -16
---
y = -2
-----
Point:: (0,-2)
----------------------
Cheers,
Stan H.